Kylah entreprises began the current amount with inventory costing $10,000, then purchased inventory at a cost of $35,000. The perpetual inventory system indicates that inventory costing $30,000 was sold during the month for $40,000. If an inventory count shows that inventory costing $14,500 is actually on hand at month-end, what amount of shrinkage occurred during the month?

Amount of Shrinkage
Answer #1

Cost of Goods available for sale = 10,000 + 35,000 = 45,000

Cost of Goods Sold = 30,000

Ending inventory = 45,000 - 30,000 = 15,000

Amount of shrinkage

= 15,000 - given ending inventory

= 15,000 - 14,500

= $500